OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 7 – 5x + 4x when x = 3 –20 4 30 18

OpenStudy (anonymous):

So the way we do this is by substitution. All you have to do is substitute x=3 into the expression. That means you replace the x in 7-5x+4x with 3. Can you try that?

OpenStudy (thread252):

4

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

But try to let the thread starter answer - it would be useful for them to solve for themself

OpenStudy (anonymous):

7 – 5x + 4x = 7 - x = 7 - 3 = 4